# Annotation Cost-Efficient Active Learning for Deep Metric Learning Driven Remote Sensing Image Retrieval

### v1

@@ Abstract
Deep metric learning (DML) has shown to be very effective for content-based image retrieval (CBIR) in remote sensing (RS). Most of DML methods for CBIR rely on a high number of annotated images to accurately learn model parameters of deep neural networks (DNNs). However, the process of gathering a large number of image annotations is time consuming and costly. To address this problem, we propose an annotation cost-efficient active learning (ANNEAL) method specifically designed for DML driven CBIR in RS. ANNEAL aims to create a small but informative training set made up of similar and dissimilar image pairs to be utilized for accurately learning a deep metric space. The informativeness of the image pairs is evaluated based on the combination of uncertainty and diversity criteria. To assess the uncertainty of image pairs, we introduce two different algorithms: 1) metric-guided uncertainty estimation (MGUE); and 2) binary classifier guided uncertainty estimation (BCGUE). MGUE algorithm automatically estimates a threshold value that acts as a ”boundary” between similar and dissimilar image pairs based on the distances in the metric space. The closer the similarity between image pairs is to the estimated threshold value the higher is their uncertainty. BCGUE algorithm automatically estimates the uncertainty of the image pairs based on the confidence of the classifier in assigning the correct similarity label. The diversity criterion is assessed through a clustering-based strategy. ANNEAL selects the most informative image pairs by combining either MGUE algorithm or BCGUE algorithm with clustering-based strategy. The selected most informative image pairs are forwarded to expert annotators to be labeled as similar or dissimilar. This way of annotating images significantly reduces the annotation cost compared to the cost of annotating images with land-use land-cover class labels. Experimental results carried out on two RS benchmark datasets demonstrate the effectiveness of our method. The code of the proposed method will be publicly available upon the acceptance of the paper.

### v2

@@ Abstract
Deep metric learning (DML) has shown to be effective for content-based image retrieval (CBIR) in remote sensing (RS). Most of DML methods for CBIR rely on a high number of annotated images to accurately learn model parameters of deep neural networks (DNNs). However, gathering such data is time-consuming and costly. To address this, we propose an annotation cost-efficient active learning (ANNEAL) method tailored to DML-driven CBIR in RS. ANNEAL aims to create a small but informative training set made up of similar and dissimilar image pairs to be utilized for accurately learning a metric space. The informativeness of image pairs is evaluated by combining uncertainty and diversity criteria. To assess the uncertainty of image pairs, we introduce two algorithms: 1) metric-guided uncertainty estimation (MGUE); and 2) binary classifier guided uncertainty estimation (BCGUE). MGUE algorithm automatically estimates a threshold value that acts as a boundary between similar and dissimilar image pairs based on the distances in the metric space. The closer the similarity between image pairs is to the estimated threshold value the higher their uncertainty. BCGUE algorithm estimates the uncertainty of the image pairs based on the confidence of the classifier in assigning correct similarity labels. The diversity criterion is assessed through a clustering-based strategy. ANNEAL combines either MGUE or BCGUE algorithm with the clustering-based strategy to select the most informative image pairs, which are then labelled by expert annotators as similar or dissimilar. This way of annotating images significantly reduces the annotation cost compared to annotating images with land-use land-cover class labels. Experimental results on two RS benchmark datasets demonstrate the effectiveness of our method. The code of this work is publicly available at https://git.tu-berlin.de/rsim/anneal_tgrs.